# What Prompts Don’t Say: Understanding and Managing Underspecification in LLM Prompts

### v1

@@ Abstract
Building LLM-powered software requires developers to communicate their requirements through natural language, but developer prompts are frequently underspecified, failing to fully capture many user-important requirements. In this paper, we present an in-depth analysis of prompt underspecification, showing that while LLMs can often (41.1%) guess unspecified requirements by default, such behavior is less robust: Underspecified prompts are 2x more likely to regress over model or prompt changes, sometimes with accuracy drops by more than 20%. We then demonstrate that simply adding more requirements to a prompt does not reliably improve performance, due to LLMs’ limited instruction-following capabilities and competing constraints, and standard prompt optimizers do not offer much help. To address this, we introduce novel requirements-aware prompt optimization mechanisms that can improve performance by 4.8% on average over baselines that naively specify everything in the prompt. Beyond prompt optimization, we envision that effectively managing prompt underspecification requires a broader process, including proactive requirements discovery, evaluation, and monitoring.
@@ 1 Introduction
As large language models (LLMs) are improving in capabilities and instruction following [30], they are increasingly integrated into business applications [e.g., 3, 7] through customized instruction prompts that can span thousands of words [e.g., 1, 2]. Conveying nuanced intentions to LLMs, however, is inherently difficult. For end users interacting directly with a model (e.g., ChatGPT), this issue is less significant, as their prompts are typically one-off and considered successful as long as they yield one satisfactory response. For developers of LLM-powered applications, the problem is much more serious, as their prompts need to generalize to many different user inputs.
@@
As an example (Figure 1), think about a developer working on an LLM-powered trip advisor application: The developer builds a prompt that specifies LLMs’ task ➀, tone ➁, and various behaviors, including avoiding transaction handling ➂ and clarify ambiguity ➃. This prompt is detailed yet still underspecified in many aspects: For example, it does not specify whether LLMs should alert users of potential weather conditions, proactively ask follow-up questions, or remind visa (entry) requirements for suggested destinations. If the LLM ends up not satisfying these requirements, it can cause frustrations and failures, such as users booking travel activities during bad weather, receiving vague recommendations, or facing denied entry due to visa issues, ultimately undermining trust in the LLM-powered applications. Indeed, failing to mention prerequisites of suggested activities has already led to a chaotic user experience with an existing LLM-powered trip advisor [6].
@@
While it is possible that developers simply do not care enough to specify these behaviors, it is equally – if not more – likely that current engineering and evaluation practices make it difficult for developers to identify such underspecification until they have already led to issues in deployment (Section 2). Even if developers do not believe these behaviors must be explicitly defined, they may expect models to act consistently along these dimensions to support more stable user mental models of LLM applications, which is, however, not the case for many unspecified requirements (Figure 1).
Figure 1: Developers often underspecify their prompts and miss user-important requirements. These unspecified requirements might be guessed by LLMs by default, but are more often inconsistent across prompts or LLMs. Such underspecification can cause frustration and failures for the end users.
@@
In this work, we characterize (Section 2) and empirically analyze (Section 3) the problem of prompt underspecification. To systematically study the effects of underspecification, we curated a set of diverse requirements across 3 representative tasks, constructed a series of prompts that specify different subsets of the requirements, and evaluated each requirement’s satisfaction rate with human-validated LLM-as-a-judge [48] on all prompts. In total, we collected 8.4k data points of LLM+Prompts’ aggregated behaviors on diverse requirements. Our analysis demonstrated that, while LLMs can indeed often (41.1%) fill in the underspecification gap, their behaviors are rather inconsistent: One version of an LLM may excel in fulfilling an unspecified requirement, but the next version can unexpectedly degrade by more than 20%. This will be a problem for continuously developing, deploying, and maintaining LLM-powered applications reliably.
@@
We then introduce requirements-aware prompt optimization as a solution strategy to deliberately communicate important requirements to the model, while leaving those already implicitly fulfilled unspecified. We show that such strategies overcome issues with existing approaches: The obvious strategy of simply specifying all requirements in the prompt does not work, due to LLMs’ limited instruction-following capabilities – their performance can drop by 19% as we specify more requirements (Section 3.5), and requirement-agnostic prompt optimization only provides limited help since they have no requirement-specific feedback. We propose and evaluate two requirement-aware prompt optimizers (Section 4): One to optimize how to specify requirements, and the other to explicitly optimize what requirements to specify. We demonstrate that both strategies work well (+4.8% accuracy), and the latter can produce shorter prompts (-43% tokens) that are easier to follow for the model.
@@
Finally, we discuss what it takes to properly manage prompt underspecification when building LLM-powered applications in practice beyond prompt optimization (Section 5): This includes proactively discovering important requirements, building reliable requirement evaluators, as well as continuously evaluating and monitoring (un-)specified requirements. We highlight the research opportunities here to support the entire process of managing prompt (under-)specification.
@@
In summary, our work makes the following contribution: (1) Characterization of the underspecification problem in LLM prompts (Section 2), (2) an empirical analysis of LLM+Prompt behaviors when underspecified (Section 3), (3) mitigation mechanisms with requirements-aware prompt optimizers (Section 4), and (4) a discussion of our vision for managing prompt underspecification (Section 5).
2 Underspecification is Amplified in LLM Prompts
While underspecification is a known challenge in traditional software engineering and machine learning pipelines [17, 11], it is significantly exacerbated in the context of LLM prompting. In this section, we articulate unique characteristics in LLM prompts that make underspecification more challenging: Unlike traditional systems, prompt engineering is informal, fast-changing, and rarely guided by systematic engineering practices [22, 43]. Moreover, LLM behaviors are highly sensitive to prompt phrasing and LLM versions, which further amplifies the effects of underspecification. A summary of comparison is in Table 1.
Table 1: Compared to traditional ML models and software, LLM prompts are more prone to underspecification, exhibit greater instability, and evolve more frequently. Prompt engineering practices and the dynamic LLM+prompt interplay both amplify these issues.
@@ 2 Underspecification is Amplified in LLM Prompts
Engineering Practices Ad-hoc, trial-and-error prompt iteration [22, 43] More structured experimentation pipelines [16] Systematic requirements engineering and design processes [40]
Artifact Evolution Frequent changes with little version control [39, 22] Periodical evolution with some version control [16] Evolution often intentionally tracked and limited [25, 8]
Behavioral Stability Highly sensitive to prompt phrasing and LLM version [9, 24] Moderate variability across training runs and datasets [11] Mostly deterministic and version-controlled
Symptoms of Underspecification Inconsistent and unexpected LLM behaviors (Section 3) Generalization failures, bias due to data underspecification [11] Software that does not meet customer needs; incorrect behavior in edge cases [36, 21]
Prompt engineering practices exacerbate underspecification. Prompts are developed with the expectation that not everything needs to be specified – ideally, LLMs should behave like a human and fill in the gaps with commonsense. This expectation encourages a “minimal-specification” prompt engineering practice: Developers begin with an initial prompt, observe violations of expected behavior, and iteratively revise through adding more instructions (i.e., specifications) [43, 22]. This trial-and-error process lacks the rigor of traditional requirements engineering [40], and exposes only a narrow slice of possible behaviors. This leaves plenty of room to miss important behaviors as developers look only at a few examples and do not necessarily know what they are looking for (Section 3.4).
Prompts and LLMs evolve much more frequently together. Prompts are not static artifacts – they are routinely modified to add new features, address failures, or adapt to updated LLMs [39]. Simultaneously, LLMs themselves undergo frequent version changes, often without clear change logs, and sometimes silently without developers’ control [24, 10]. This dual axis of prompt and model changes introduces behavioral drift in addition to traditional requirements changes. In contrast, traditional software change is often carefully managed [25, 8] and most ML pipelines evolve more slowly.
LLM behaviors are highly unstable and sensitive to (under-)specification. LLM behaviors are known to be unstable and sensitive to prompt wordings [9]. Similarly, given the same prompt, different LLMs exhibit different behaviors [13], even just across small version updates [10]. As we will show later (Section 3), LLM behaviors are especially unstable in underspecified settings, where LLM behaviors depend on implicit assumptions that are not guaranteed to transfer across changes.
@@ 3 How do LLM+Prompts Behave when Underspecified?
While LLM+Prompts are generally known to be unstable, existing work mostly studies their stability on specified requirements in the prompts. To quantitatively measure LLM+Prompts’ behaviors on unspecified requirements, we design an experiment as follows: We collect a set of 60 plausible requirements across 3 tasks and create validated, automated validators for each requirement. We then create prompts with subsets of the requirements and measure how well the model’s outputs meet the (un-)specified requirements using their validators.
Our analysis starts with showing that LLMs can often guess unspecified requirements, but these behaviors are inconsistent (Section 3.2) and more likely to degrade with model updates (Section 3.3). We next analyze why identifying unspecified requirements is fundamentally hard with current practices, especially when requirements are conditional or rarely violated (Section 3.4). Finally, we show that an obvious solution of specifying all requirements in a single prompt actually hurts performance, due to LLMs’ limited ability to follow long, complex instructions (Section 3.5).
@@ 3.1 Experiment Setups
Tasks and data. We selected three tasks based on Anthropic’s report measuring AI usage patterns [14], covering diverse tasks that can be integrated in commercial applications, from software engineering (code-explain), travel industry (trip-advisory), to E-commerce business (product-gen).
We re-purposed three existing datasets to run the LLM+Prompts on: Commitpackft [27] for code explanation, a subset of UltraChat [12] for trip advisory, and Amazon ESCI [32] for product description generation. From each dataset, we sampled 200 examples, and split them into training, validation, and test data with 15/35/50 split. More details on task and data can be found at Appendix A.1– A.2.
Requirements. The key setup of our experiment is to curate a list of plausible requirements for each task we study. [^1] Note that there is no fixed set of requirements for a task, as different developers or users may have different priorities. We intentionally curate requirements that are plausible and likely shared. We consider three different sources to curate the requirements:
@@
  • LLM-supported brainstorming. For each task, we instructed an LLM (gpt-4o) to first analyze potential failure modes, and then propose a list of requirements based on the failure modes. This simulates expert-driven requirements elicitation methods, generating comprehensive requirements by anticipating potential system failures [4, 42].
  • LLM-supported error analysis. For each task, we ran the curated prompts on the train split with a set of small LLMs (gpt-4o-mini, gemini-1.5-flash, llama3.2-11b). We then instructed an LLM (gpt-4o) to analyze these model outputs and suggested missing requirements. This simulates error analysis [28] and produces requirements that are grounded in real model mistakes.
@@
Different from existing instruction-following benchmarks [31, 50], this setup allows us to cover more real-world-based, diverse requirements (vs. synthetic ones), and much larger complex prompts.
Requirement validators. For each model output, we validate how well they satisfy each (specified or unspecified) requirement, either using a Python script or LLM-as-a-judge [48]. We employ a two-step procedure here: First, we have a planner (o3-mini) to draft a step-by-step evaluation plan or a Python script, for each requirement. Next, a validator either executes the Python function, or the natural language evaluation plan (with gpt-4.1-mini) on each example to produce the final judgment. We iterate the validators for each requirement on the train split to achieve high human-LLM agreement. We manually validate the LLM validators with a sampled test subset, achieving an average of 95.6% human-LLM agreement. More details of our human validation can be found in Appendix A.6.
@@ 3.2 LLMs are often able to guess unspecified requirements but lack stability
Setups. We study three models, one smaller open-sourced model Llama-3.3-70B-Instruct, one more powerful closed-sourced model gpt-4o-2024-08-06, and one reasoning model o3-mini. We calculate the average accuracy and standard deviation for each requirement when specified or unspecified, and compare their distribution.
Figure 2: While LLMs perform worse (-22.6% on average) when a requirement is unspecified (top), they are often (41.1% on average) able to guess unspecified requirements ($\geq 0.98$ accuracy), with increased capabilities (bottom).
Results. Unsurprisingly, we generally observe that LLM+Prompts are less likely to implement a requirement when unspecified (-22.6% on average, up to -93.1%, Figure 2). However, we also found that LLMs are often able to guess unspecified requirements – in 41.1% of cases, they are able to achieve more than 98% accuracy on unspecified requirements. Indeed, for 65% of all requirements, we found at least one LLM+prompt combination that is able to guess it without explicit specification.
Comparing across models, we found that stronger LLMs are more likely to guess requirements: o3-mini can guess 44.7% of unspecified requirements, a 20.2% increase compared to Llama3.3-70b-instruct, possibly benefiting from its capabilities to “infer” requirements in reasoning.
@@
While LLMs are often able to guess unspecified requirements, we found them also to be generally less robust with unspecified requirements across different prompts: Different prompts can guess unspecified requirements completely differently. On average, they have a standard deviation of 8.9%, a more than 2x increase compared to when they are specified.
Implications. While LLMs do follow many unspecified requirements, we found there is no clear pattern when they do – developers also do not seem to identify ones that need to be specified. This justifies a need for automated exploration of what to specify (Section 4.2) and properly managing and testing all relevant requirements, whether specified in the prompt or not (Section 5).

### v2

@@ Abstract
Prompt underspecification is a common challenge when interacting with LLMs. In this paper, we present an in-depth analysis of this problem, showing that while LLMs can often infer unspecified requirements by default (41.1%), such behavior is fragile: Under-specified prompts are 2x as likely to regress across model or prompt changes, sometimes with accuracy drops exceeding 20%. [^1] Code and data shared in https://github.com/malusamayo/underspec-analysis. This instability makes it difficult to reliably build LLM applications. Moreover, simply specifying all requirements does not consistently help, as models have limited instruction-following ability and requirements can conflict. Standard prompt optimizers likewise provide little benefit. To address these issues, we propose requirements-aware prompt optimization mechanisms that improve performance by 4.8% on average over baselines. We further advocate for a systematic process of proactive requirements discovery, evaluation, and monitoring to better manage prompt underspecification in practice.
What Prompts Don’t Say: Understanding and Managing Underspecification in LLM Prompts
Chenyang Yang  Yike Shi  Qianou Ma  Michael Xieyang Liuthanks: Now at Google Deepmind. Christian KästnerTongshuang Wu Carnegie Mellon University
@@ 1 Introduction
As large language models (LLMs) are improving in capabilities and instruction following Ouyang et al. (2022), they are increasingly integrated into commercial applications through customized instruction prompts that can span thousands of words (e.g., All-Hands-AI, 2025; Anthropic, 2025). Conveying nuanced intentions to LLMs, however, is inherently difficult. This issue is less significant for end users, as their prompts are typically one-off and considered successful as long as they yield one satisfactory response throughout their interactions. For LLM application developers, the problem is much more serious, as their prompts need to generalize to many different usage scenarios.
@@
As an example (Figure 1), consider a developer building an LLM-powered trip advisor: The detailed prompt may specify the task ➀, tone ➁, and certain behaviors such as avoiding transactions ➂ and clarifying ambiguity ➃, yet it remains underspecified in many aspects: For example, it does not specify whether LLMs should warn about weather, proactively ask follow-up questions, or remind users of visa (entry) requirements. If the LLM ends up not satisfying these requirements, it can cause frustrations and failures, such as users booking activities during bad weather, receiving vague recommendations, or facing denied entry due to visa issues, ultimately undermining trust in the LLM-powered applications. Indeed, failing to mention prerequisites of suggested activities has already caused chaotic experiences in an existing LLM trip advisor Bernal and Hoover (2024).
@@
We define underspecification as the omission of essential requirements in a prompt, such that multiple valid but inconsistent behaviors remain possible. While it is possible that developers simply do not care enough to specify these behaviors, it is equally – if not more – likely that current engineering and evaluation practices make it difficult for developers to identify such underspecification until they have already led to issues in deployment (Section 2). Even if developers do not believe these behaviors must be explicitly defined, they may expect models to act consistently along these dimensions to support more stable user mental models of LLM applications, which is, however, not the case for many unspecified requirements (Figure 1).
Figure 1: Developers often underspecify prompts and miss user-important requirements, leading to divergent behaviors. We analyze the challenges underspecification presents and propose mechanisms to manage the problem.
@@
In this work, we characterize (Section 2) and empirically analyze (Section 3) the problem of prompt underspecification. To systematically study the problem, we curated a set of diverse requirements across 3 representative tasks, constructed a series of prompts that specify different subsets of the requirements, and evaluated each requirement’s satisfaction rate with human-validated LLM-as-a-judge Zheng et al. (2023). In total, we collected 8.4k data points of LLM+Prompts’ aggregated behaviors on diverse requirements. Our analysis demonstrated that, while LLMs can indeed often (41.1%) fill in the underspecification gap, their behaviors are rather inconsistent: One version of an LLM may excel in fulfilling an unspecified requirement, but the next version can unexpectedly degrade by more than 20%. This will be a problem for continuously developing, deploying, and maintaining LLM-powered applications reliably.
@@
We then introduce requirements-aware prompt optimization as a solution strategy to deliberately communicate important requirements to the model, while leaving those already implicitly fulfilled unspecified (Section 4). We show that such strategies overcome issues with existing approaches: The obvious strategy of simply specifying all requirements in the prompt does not work, due to LLMs’ limited instruction-following capabilities – their performance can drop by 19% as we specify more requirements (Section 3.4), and requirement-agnostic prompt optimization only provides limited help since they have no requirement-specific feedback. We propose and evaluate two requirement-aware prompt optimizers: One to optimize how to specify requirements, and the other to explicitly optimize what requirements to specify. We demonstrate that both strategies work well (+4.8% accuracy), and the latter can produce shorter prompts (-43% tokens) that are easier to follow for the model.
@@
Finally, we discuss how to manage prompt underspecification when building LLM applications in practice beyond prompt optimization (Section 5): This includes proactively discovering important requirements, building reliable requirement evaluators, as well as continuously evaluating and monitoring (un-)specified requirements. We highlight the research opportunities here to support the entire process of managing prompt (under-)specification.
@@
In summary, our work makes the following contribution: (1) Characterization of the prompt underspecification problem (Section 2), (2) an empirical analysis of LLM+Prompt behaviors when underspecified (Section 3), (3) mitigation mechanisms with requirements-aware prompt optimizers (Section 4), and (4) a discussion of our vision for managing prompt underspecification (Section 5).
2 Underspecification is Amplified in Prompting LLMs
While underspecification is a known challenge in traditional software engineering and machine learning pipelines Kästner et al. (2021); D’Amour et al. (2022), it is significantly amplified in prompting LLMs. We next elaborate on two key factors behind this amplification: (a) the lack of rigor in prompt engineering practices, and (b) the rapid co-evolution of prompts and LLMs. A summary of these differences is provided in Table 1.
@@ 2 Underspecification is Amplified in Prompting LLMs
Engineering Practices Ad-hoc, trial-and-error prompt iteration Liang et al. (2024); Zamfirescu-Pereira et al. (2023) More structured experimentation pipelines Huyen (2022) Systematic requirements engineering and design processes Van Lamsweerde (2009)
Artifact Evolution Frequent changes with little version control Tafreshipour et al. (2024); Liang et al. (2024) Periodical evolution with some version control Huyen (2022) Evolution often intentionally tracked and limited McConnell (1998); Bogart et al. (2021)
Consequences Inconsistent and unexpected LLM behaviors (Section 3) Generalization failures, model biases D’Amour et al. (2022) Software that does not meet customer needs; incorrect behavior in edge cases Sommerville (2015)
Table 1: Compared to traditional ML models and software, LLM prompts are more prone to underspecification, less stable, and evolve more frequently. Prompt engineering practices and further amplify these issues.
Ad-hoc prompt engineering amplifies underspecification. Prompts are developed with the expectation that not everything needs to be specified – ideally, LLMs should behave like a human and fill in the gaps with commonsense. This expectation encourages a “minimal-specification” prompt engineering practice: Developers begin with an initial prompt, observe violations of expected behavior, and iteratively revise through adding more instructions (i.e., specifications) Zamfirescu-Pereira et al. (2023); Liang et al. (2024). This trial-and-error process lacks the rigor of traditional requirements engineering Van Lamsweerde (2009), and exposes only a narrow slice of possible behaviors.
Yet many requirements demand more systematic discovery efforts: Some are conditional (e.g., a “mentioning visa” requirement applies only to international travel). Others are infrequently violated yet critical (e.g., failing “no dangerous activities suggested” is high-stake even uncommon). Still others are difficult to recognize (e.g., verifying “suggested sites are geographically close”). Because such requirements are easily overlooked in ad-hoc iteration, underspecification persists even as prompts appear to perform well on inspected examples.
Rapid Prompt-LLM co-evolution amplifies underspecification. Prompts are not static artifacts – they are routinely modified to add new features, address failures, or adapt to updated LLMs Tafreshipour et al. (2024). Because their specifications are written in natural language, such revisions are easy and quick to make, encouraging frequent, informal changes. At the same time, LLMs themselves evolve rapidly, often without clear change logs, and sometimes silently without developers’ control Ma et al. (2024).
This dual, fast-moving evolution of both prompts and models introduces continual behavioral drift: Previously validated behaviors may no longer hold, requiring repeated rediscovery of requirements. The result is a growing maintenance burden for developers, who must continually re-evaluate prompt behavior to keep up with both axes of change. In contrast, traditional software change is often carefully managed McConnell (1998); Bogart et al. (2021), and most ML pipelines evolve more slowly.
@@ 3 How do LLM+Prompts Behave when Underspecified?
The difficulty of reliably discovering unspecified requirements raises a natural follow-up question: What happens when such requirements are left out of the prompt? While LLM+Prompts are generally known to be unstable, existing work mostly studies their stability on specified requirements. To quantitatively measure LLM+Prompts’ behaviors on unspecified requirements, we design an experiment as follows: We collect a set of 60 plausible requirements across 3 tasks and create human-validated, automated validators for each requirement. We then create prompts with subsets of the requirements and measure how well the model’s outputs meet the (un-)specified requirements with validators.
Our analysis starts with showing that LLMs can often guess unspecified requirements, but these behaviors are inconsistent (Section 3.2) and more likely to degrade with model updates (Section 3.3). We show that an obvious solution of specifying all requirements in a single prompt actually hurts performance, due to LLMs’ limited ability to follow long, complex instructions (Section 3.4).
@@ 3.1 Experiment Setups
Tasks and data. We selected three tasks based on Anthropic’s report measuring AI usage patterns Handa et al. (2025), covering diverse tasks that can be integrated in commercial applications, from software engineering (code-explain), travel industry (trip-advisory), to E-commerce business (product-gen).
We re-purposed three existing datasets to run the LLM+Prompts on: Commitpackft Muennighoff et al. (2023) for code explanation, a subset of UltraChat Ding et al. (2023) for trip advisory, and Amazon ESCI Reddy et al. (2022) for product description generation. From each dataset, we sampled 200 examples, and split them into training, validation, and test data with 15/35/50 split. More details can be found at Appendix A.1– A.2.
Requirements. The key setup of our experiment is to curate a list of plausible requirements for each task we study. [^2] Note that there is no fixed set of requirements for a task, as different developers may have different priorities. We aim to curate requirements that are plausible and likely shared. We consider three different sources to curate the requirements:
@@
  • LLM-supported brainstorming. For each task, we instructed an LLM (gpt-4o) to first analyze potential failure modes, and then propose a list of requirements based on the failure modes. This simulates expert-driven requirements elicitation methods, generating comprehensive requirements by anticipating potential system failures Barzamini et al. (2022); Yang et al. (2023).
  • LLM-supported error analysis. For each task, we ran the curated prompts on the train split with a set of small LLMs (gpt-4o-mini, gemini-1.5-flash, llama3.2-11b). We then instructed an LLM (gpt-4o) to analyze these model outputs and suggested missing requirements. This simulates error analysis to produce requirements grounded in real model mistakes.
@@
Requirement validators. For each model output, we validate how well they satisfy each (specified or unspecified) requirement, either using a Python script or an LLM validator Zheng et al. (2023). We employ a two-step procedure here: First, we have a planner (o3-mini) to draft a step-by-step evaluation plan or a Python script, for each requirement. Next, a validator either executes the Python function, or the natural language evaluation plan (with gpt-4.1-mini) on each example to produce the final judgment. We iterate the validators for each requirement on the train split to achieve high human-LLM agreement. We manually validate the LLM validators with a sampled test subset, achieving an average of 95.6% human-LLM agreement. More details of our human validation can be found in Appendix A.6.
@@ 3.2 LLMs are often able to guess unspecified requirements but lack stability
Setups. We study three models, one smaller open-sourced model Llama-3.3-70B-Instruct, one big closed-sourced model gpt-4o-2024-08-06, and one reasoning model o3-mini. We calculate the average accuracy and standard deviation for each requirement when specified or unspecified, and compare their distribution.
Results. Unsurprisingly, we generally observe that LLM+Prompts are less likely to implement a requirement when unspecified – accuracy drops by an average of 22.6% (and up to 93.1%) compared to when the requirement is explicitly stated (Figure 2). Yet, not all requirements are equally affected. We found that LLMs are often able to guess unspecified requirements – in 41.1% of cases, they are able to achieve more than 98% accuracy on unspecified requirements. Indeed, for 65% of all requirements, we found at least one LLM+prompt combination that is able to guess it without explicit specification.
Comparing across models, stronger LLMs are more likely to guess requirements: o3-mini can guess 44.7% of unspecified requirements, a 20.2% increase compared to Llama3.3-70b-instruct, possibly due to extra test-time compute to “infer” requirements in reasoning.
Figure 2: While LLMs perform worse (-22.6% avg.) when a requirement is unspecified (top), they are often (41.1% avg.) able to guess unspecified requirements ($\geq 0.98$ accuracy), with increased capabilities (bottom).
@@
While LLMs are often able to guess unspecified requirements, we found them less robust with unspecified requirements across different prompts: Different prompts can guess unspecified requirements completely differently. On average, they have a standard deviation of 8.9%, a more than 2x increase compared to when they are specified. We observe even a stronger effect when explicitly controlling for requirement conflicts (7.9% unspecified vs. 0.8% specification, details in Appendix C.3).
Implications. While LLMs do follow many unspecified requirements, their success depends on how a specific model is post-trained. Developers seem to struggle with identifying ones that need to be specified. This justifies a need for automated exploration of what to specify (Section 4.2) and properly managing and testing all relevant requirements, whether specified in the prompt or not (Section 5).